# GraphTeam: Facilitating Large Language Model-based Graph Reasoning via Multi-Agent Collaboration

### v4

@@ 4.6 Hyper-parameter Analysis
As shown in Figure 6, we conduct experiments to study the influence of key hyper-parameters using the NLGraph Wang et al. (2024a) benchmark. For the number of candidate experiences $N_{exp}$ in knowledge base construction, our experiments show that larger $N_{exp}$ values correlate with better system performance, as they provide more candidate experiences for problem-solving assistance. The similarity matching threshold $\delta$ between experiences and documents achieves optimal problem-solving performance at $0.85$, enabling correct matching of similar experiences when available and document retrieval when experiences are absent. Increasing the maximum retries of the coding agent and the number of self-checking iterations of the answer agent improves system performance, but with diminishing returns as these values increase, eventually plateauing. These findings provide valuable insights for optimizing the system’s configuration and balancing performance gains against computational costs.
@@ A.1 Problem-Solving Experience Collection
[…]
@@
2:     if $|\mathcal{K}_{exp}^{t}|$ $<$ $N_{exp}$ then
[…]
@@
[…]
@@
[…]
@@ B.1 Dataset Statistics
Task Category Output Format Graph Size Total Experiences Test Set Size
NLGraph Macro-level, Micro-level, GNN yes/no, digits, list/set, others $\sim 10^{1}$ 80 1,000
Talk like a Graph Basic, Macro-level yes/no, digits, list/set $\sim 10^{1}$ 80 8,000
GraphInstruct Basic, Macro-level, Micro-level yes/no, digits, list/set $\sim 10^{1}$ 170 5,100
GraphWiz Macro-level, Micro-level yes/no, digits, list/set $\sim 10^{2}$ 90 3,600
LLM4DyG Basic, Micro-level yes/no, digits, list/set $\sim 10^{1}$ 90 2,700
GNN-AutoGL GNN others $\sim 10^{3}$ 10 200
@@ B.2 GNN-AutoGL Benchmark Construction
Given the cora dataset, can you use autogl to train a $2$-layer $gcn$ model with hidden dimension as $[128,256]$, dropout as $0.6$ and predict the accuracy ?
@@ B.3 Metric Details
Specifically, we extract the generated codes from model responses, and compare the execution results with the reference solutions to assess whether the model-generated results match the target answers. For the first five benchmarks Fatemi et al. [2024], Chen et al. [2024a], Zhang et al. [2024b], Luo et al. [2024], Wang et al. [2024a], the evaluation scores are binary based on the exact matches between execution results and target answers. For GNN-AutoGL, we first evaluate whether the code can be executed without errors; if it cannot, it scores 0. Otherwise, we check whether the hyper-parameters (e.g., model name, dataset name, hidden dimension, dropout) meet the requirements. Each correct hyper-parameter earns partial credit, and if all hyper-parameters are correct, the score is 1.
@@ Appendix C Additional Experiments
Category Models NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
@@ C.4.1 NLGraph
Connect Cycle Topo. Sort Shortest Path Max. Flow Bipartite Graph Hamilton Path GNN Overall Average
@@
As shown in Table 5, we present detailed results across subclasses of NLGraph. GraphTeam (GPT-4o-mini) achieves perfect scores in multiple categories (including Cycle, Maximum Flow, Bipartite Graph, and Hamilton Path), with its lowest performance being 94.8 in Topology Sort. These results substantially exceed the previous SOTA Ouyang et al. [2024] and GPT-4o-mini baseline.
@@ C.4.2 Talk Like a Graph
Node Count Edge Count Edge Existence Node Degree Connected Nodes Cycle Check Shortest Path Triangle Counting Overall Average
@@ C.4.3 GraphInstruct
Neighbor Bipatite Edge Connectivity Degree DFS Predecessor Topological Sort Connected Component
@@
Common Neighbor Hamiltonian Path Jaccard Shortest Path Diameter PageRank Maximum Flow MST Overall Average
@@ C.4.4 GraphWiz
Cycle Connect Bipartite Topology Shortest Triangle Flow Hamilton Subgraph Overall Average
@@
The results are shown in Table 8. Although both GPT-4o-mini and GraphTeam (GPT-4o-mini) can solve the Hamilton problem conceptually, their performance is substantially impacted by the inability to match the complex output format required by the evaluation script of GraphWiz. The previous SOTA model Chen et al. [2024a] is fine-tuned on extensive datasets, and thus able to meet the required output format. Nevertheless, GraphTeam achieves an overall accuracy of 88.1% on GraphWiz, demonstrating strong graph reasoning capabilities.
@@ C.4.5 LLM4DyG
When Link When Connect When Tclosure Neighbor at Time Neighbor in Periods Check Tclosure Check Tpath Find Tpath Sort Edge Overall Average

### v5

@@ 4.6 Hyper-parameter Analysis
As shown in Figure LABEL:fig:hyper-parameter, we conduct experiments to study the influence of key hyper-parameters using the NLGraph [42] benchmark. For the number of candidate experiences $N_{exp}$ in knowledge base construction, our experiments show that larger $N_{exp}$ values correlate with better system performance, as they provide more candidate experiences for problem-solving assistance. The similarity matching threshold $\delta$ between experiences and documents achieves optimal problem-solving performance at $0.85$, enabling correct matching of similar experiences when available and document retrieval when experiences are absent. Increasing the maximum retries of the coding agent and the number of self-checking iterations of the answer agent improves system performance, but with diminishing returns as these values increase, eventually plateauing. These findings provide valuable insights for optimizing the system’s configuration and balancing performance gains against computational costs.
@@ A.1 Problem-Solving Experience Collection
The pseudo code for building the experience base is shown in Alg. 2. Each problem in the training or validation datasets is represented as a triple $(q,a,t)$ of question, answer and type. The solver is formalized as function $\operatorname{Solver}(\operatorname{Question}(q),e)$ with the output of question agent and an optional experience entry $e$ as inputs, and can be seen as a simplified version of GraphTeam without the entire experience base. Besides generated answer, the outputs of the solver also include meta information such as codes and thinking processes. If the solver can correctly answer a problem, we will add a triple $(q,a,\text{meta\_info})$ of question, answer and meta information into the candidate experiences $\mathcal{K}_{exp}^{t}$. Then for each experience $e\in\mathcal{K}_{exp}^{t}$, we evaluate its utility by solving the problems of the same type $t$ in the validation set $\mathcal{D}_{\text{valid}}$ with experience $e$ as augmentation. Then the experience base $\mathcal{K}_{exp}$ collects the experience with the highest utility in each $\mathcal{K}_{exp}^{t}$.
@@
2:  if $|\mathcal{K}_{exp}^{t}|$ << $N_{exp}$ then
3:   $(\text{answer},\text{meta\_info})\leftarrow\text{Solver}(\operatorname{Question}(q))$
@@
5:    $\mathcal{K}_{exp}^{t}\leftarrow\mathcal{K}_{exp}^{t}\cup\{(q,a,\text{meta\_info})\}$
@@
11:   $(\text{answer},\text{meta\_info})\leftarrow\text{Solver}(\operatorname{Question}(q),e)$
@@ B.1 Dataset Statistics
Task Category
Output Format
Graph Size
Total Experiences
Test Set Size
NLGraph
Macro-level, Micro-level, GNN
yes/no, digits, list/set, others
$\sim 10^{1}$
80
1,000
Talk like a Graph
Basic, Macro-level
yes/no, digits, list/set
$\sim 10^{1}$
80
8,000
GraphInstruct
Basic, Macro-level, Micro-level
yes/no, digits, list/set
$\sim 10^{1}$
170
5,100
GraphWiz
Macro-level, Micro-level
yes/no, digits, list/set
$\sim 10^{2}$
90
3,600
LLM4DyG
Basic, Micro-level
yes/no, digits, list/set
$\sim 10^{1}$
90
2,700
GNN-AutoGL
GNN
others
$\sim 10^{3}$
10
200
@@ B.2 GNN-AutoGL Benchmark Construction
Given the cora dataset, can you use autogl to train a 22-layer gcngcn model with hidden dimension as $[128,256]$, dropout as $0.6$ and predict the accuracy ?
@@ B.3 Metric Details
Specifically, we extract the generated codes from model responses, and compare the execution results with the reference solutions to assess whether the model-generated results match the target answers. For the first five benchmarks [12, 7, 49, 30, 42], the evaluation scores are binary based on the exact matches between execution results and target answers. For GNN-AutoGL, we first evaluate whether the code can be executed without errors; if it cannot, it scores 0. Otherwise, we check whether the hyper-parameters (e.g., model name, dataset name, hidden dimension, dropout) meet the requirements. Each correct hyper-parameter earns partial credit, and if all hyper-parameters are correct, the score is 1.
@@ Appendix C Additional Experiments
Category Models NLGraph
Talk like a Graph
GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
@@ C.4.1 NLGraph
Connect
Cycle
Topo. Sort
Shortest Path
Max. Flow
Bipartite Graph
Hamilton Path
GNN
Overall
Average
@@
As shown in Table 5, we present detailed results across subclasses of NLGraph. GraphTeam (GPT-4o-mini) achieves perfect scores in multiple categories (including Cycle, Maximum Flow, Bipartite Graph, and Hamilton Path), with its lowest performance being 94.8 in Topology Sort. These results substantially exceed the previous SOTA [32] and GPT-4o-mini baseline.
@@ C.4.2 Talk Like a Graph
Node Count
Edge Count
Edge Existence
Node Degree
Connected Nodes
Cycle Check
Shortest Path
Triangle Counting
Overall Average
@@ C.4.3 GraphInstruct
Neighbor
Bipatite
Edge
Connectivity
Degree
DFS
Predecessor
Topological
Sort
Connected
Component
@@
Common
Neighbor
Hamiltonian
Path
Jaccard
Shortest
Path
Diameter
PageRank
Maximum
Flow
MST
Overall
Average
@@ C.4.4 GraphWiz
Cycle
Connect
Bipartite
Topology
Shortest
Triangle
Flow
Hamilton
Subgraph
Overall
Average
@@
The results are shown in Table 8. Although both GPT-4o-mini and GraphTeam (GPT-4o-mini) can solve the Hamilton problem conceptually, their performance is substantially impacted by the inability to match the complex output format required by the evaluation script of GraphWiz. The previous SOTA model [7] is fine-tuned on extensive datasets, and thus able to meet the required output format. Nevertheless, GraphTeam achieves an overall accuracy of 88.1% on GraphWiz, demonstrating strong graph reasoning capabilities.
@@ C.4.5 LLM4DyG
When
Link
When
Connect
When
Tclosure
Neighbor
at Time
Neighbor
in Periods
Check
Tclosure
Check
Tpath
Find
Tpath
Sort
Edge
Overall
Average